A new labor contract provides for a wage increase of $1 per hour and a reduction of 5 hours in the workweek. A worker who received $500 per week under the old contract will receive $495 per week under the new contract. How long was the workweek under the old contract?
Answer by jorel555(1290)

This question involves two variables; Hours and wages. Thus
HW=500
W=500/H
Under the new contract:
(H-5)(W+1)=495
(H-5)((500/H)+1)=495
500-2500/H+H-5=495
-2500/H+H=0
H^2-2500=0
H^2=2500
H=50
The old workweek was 50 hours long!!!!!!!!